OpenStudy (tia_sweetheart17):

solve for y: -4+6=-6y-6

OpenStudy (ranga):

-4 + 6 = -6y - 6 simplify the left 2 = -6y - 6 Isolate the y. First add 6 to both sides: 2 + 6 = -6y 8 = -6y divide both sides by -6 8/(-6) = y y = -8/6 simplify y = -4/3